### v2

# Beyond Prompt Content: Enhancing LLM Performance via Content-Format Integrated Prompt Optimization

Abstract
Large Language Models (LLMs) have shown significant capability across various tasks, with their real-world effectiveness often driven by prompt design. While recent research has focused on optimizing prompt content, the role of prompt formatting—a critical but often overlooked dimension—has received limited systematic investigation. In this paper, we introduce Content-Format Integrated Prompt Optimization (CFPO), an innovative methodology that jointly optimizes both prompt content and formatting through an iterative refinement process. CFPO leverages natural language mutations to explore content variations and employs a dynamic format exploration strategy that systematically evaluates diverse format options. Our extensive evaluations across multiple tasks and open-source LLMs demonstrate that CFPO demonstrates measurable performance improvements compared to content-only optimization methods. This highlights the importance of integrated content-format optimization and offers a practical, model-agnostic approach to enhancing LLM performance. Code is available at https://github.com/HenryLau7/CFPO.
Beyond Prompt Content: Enhancing LLM Performance via
Content-Format Integrated Prompt Optimization
Yuanye Liu∗†   Jiahang Xu∗⋄   Li Lyna Zhang   Qi Chen   Xuan Feng Yang ChenZhongxin GuoYuqing YangPeng Cheng Fudan University    Microsoft Research Asia
\$*\$\$*\$footnotetext: Equal contribution.$\dagger$\$$\dagger$\$footnotetext: Yuanye Liu did the work during an internship at MSRA.$\diamond$\$$\diamond$\$footnotetext: Corresponding author: jiahangxu@microsoft.com
1 Introduction
Figure 1: The crucial role of prompt formatting and its interaction with content. (A): Model-specific format biases: Illustrates the performance sensitivity of two LLMs to different format styles on the GSM8K task, showing substantial variability in the effectiveness of 10 randomly selected formats. (B): For seven different prompt contents evaluated across 24 distinct formats, performance variations show the complex, interdependent relationship between prompt content and structure, demonstrating that no single format universally maximizes effectiveness.
Figure 2: Illustration of the CFPO pipeline within a single iteration round. In the initial Component-wise Content Optimization stage, case-diagnosis and Monte-Carlo sampling are employed for content mutation. Subsequently, the Format Optimization stage identifies the most suitable format for each content candidate. The yellow dashed line indicates where the LLM optimizer is employed to guide the optimization process.
Large Language Models (LLMs) have demonstrated impressive achievements across various domains OpenAI (2024a). The effectiveness of LLMs in real-world applications is fundamentally dependent on the design of effective prompts, which serve as an essential interface between human users or developers and the LLM system. Studies have shown that expert-designed prompts could significantly enhance LLM performance (Brown et al., 2020; Wei et al., 2023; Schulhoff et al., 2024).
However, manual design of prompts presents significant challenges, primarily due to the high sensitivity of LLMs to subtle variations in prompt characteristics, including both textual content and structural format (Jiang et al., 2022; Zamfirescu-Pereira et al., 2023; Salinas and Morstatter, 2024). These sensitivities are further complicated by variations across different models and tasks Zhuo et al. (2024); Sclar et al. (2024). To alleviate these difficulties, automated prompt optimization techniques, often leveraging the power of LLMs themselves, have proven to be an effective approach to adapt and refine prompts Pryzant et al. (2023); Schnabel and Neville (2024); Yang et al. (2024). However, existing research primarily focuses on optimizing prompt content, while overlooking a critical and largely unexplored dimension: the prompt formatting.
Our preliminary investigations, as illustrated in Figure 1, provide valuable insights into the role of prompt format in prompt optimization. We have observed that different LLMs display distinct preferences, with some formats performing well on one model but failing on another. This suggests sophisticated, model-specific format biases Sclar et al. (2024). Furthermore, we have identified a complex interplay between prompt content and format, where no single format consistently outperforms others across all contents. This lack of a universally optimal format highlights the impracticality of predefining format, and underscores the need for a joint optimization approach that treats prompt content and format as interdependent variables.
Figure 3: An illustrative example of our Structured Prompt Template. This template systematically organizes the prompt into distinct components, each serving a specific functional role. When formulating a prompt, the template first employs a Query format to present examples and queries, and then integrates all content components via the Prompt Renderer to construct the comprehensive prompt string.
To address these limitations, we introduce Content-Format Integrated Prompt Optimization (CFPO), an innovative methodology that concurrently optimizes both prompt content and format through an iterative refinement process. CFPO employs distinct optimization strategies tailored to the unique search spaces of content and format. Content optimization is guided by performance feedback and Monte Carlo sampling, leveraging natural language mutations to enhance prompt effectiveness. For format optimization, CFPO explores a discrete set of format options through a dynamic exploration strategy designed to identify optimal formats without requiring a prior knowledge.
Specifically, CFPO’s format optimizer leverages the principles of structured thinking, operating along two key dimensions: the Prompt Renderer, which governs the organizational structure of all components within a prompt (He et al., 2024), and the Query Format, which dictates the presentation of in-context learning examples and queries (Voronov et al., 2024a; Salinas and Morstatter, 2024). By integrating these two dimensions, CFPO defines a structured template that effectively distinguishes between content and format types, enabling the efficient identification of high-performing prompts.
Our primary contributions are threefold: (1) We propose CFPO, an innovative approach to simultaneously optimizes prompt content and format using an iterative process. (2) We introduce an efficient strategy for dynamic format optimization that generates new formats in an iterative manner and evaluates formats instance through a scoring system to select the best option. (3) Through extensive evaluations across diverse tasks and multiple open-source LLMs, we demonstrate that CFPO consistently improves LLM performance in a measurable and effective manner.
2 Related Work
Optimization via LLM The remarkable capacity of LLMs has been demonstrated in various tasks as optimizers, leveraging their ability to enhance performance, such as code generation (Haluptzok et al., 2023; Zelikman et al., 2024; Askari et al., 2024), tool-making (Cai et al., 2024), and agent system design (Hu et al., 2024). However, recent studies indicate that LLMs face significant challenges in achieving completely automatic optimization. These models often rely on human intervention for designing workflows and struggle with tasks requiring complex decomposition and iterative refinement (Zhang et al., 2024; Li et al., 2024).
Automatic Prompt Optimization Automatic prompt optimization plays a crucial role in enhancing the performance of LLMs by refining prompts without requiring human intervention. Various approaches have been explored to search for the optimal prompt, including reinforcement learning (Zhang et al., 2023), Monte Carlo Search (Zhou et al., 2023), Monte Carlo Tree Search (MCTS) (Wang et al., 2024b), feedback-based methods (Pryzant et al., 2023; Das et al., 2024), and agent-driven frameworks (Wang et al., 2024a; Khattab et al., 2024; WHO, 2023). While these methods focus on optimizing the overall prompt, they often lack the capability for fine-grained modifications. (Khattab et al., 2024; Schnabel and Neville, 2024) introduce phrase-level mutations, but they fail to address format mutations or implement them in a systematic manner.
Prompt structure and format Structured prompting, which organizes prompts into distinct components such as instructions, examples, and queries, holds significant potential in prompt engineering (Fernando et al., 2023). Empirical rules for prompt design always lack integration with automatic optimization techniques, limiting their scalability and effectiveness (Nigh, 2023; Google, 2024). Frameworks like LangGPT (Wang et al., 2024a) have introduced structured prompting paradigms, emphasizing reusable designs inspired by programming principles. However, these efforts primarily focus on content-level refinements and fail to adequately address the critical role of prompt formatting. Studies have highlighted the impact of formatting on prompt performance (Salinas and Morstatter, 2024). Sclar et al. (2024) revealed that modifications to separators and spacing within a query could substantially impact performance. He et al. (2024) reveals that the format of prompts significantly impacts GPT-based models’ performance, with no single format excelling universally. Voronov et al. (2024b) focus on the format of few-shot examples and suggests that it is beneficial to maintain a consistent format across examples. However, despite the recognition of formatting’s importance, there remains a lack of comprehensive understanding regarding the optimization of prompt format in a systematic manner.
Figure 4: Built-in formats and rendering effects in our initial format pool. The final format configuration is achieved by selecting and combining elements from both the Prompt Renderer and the Query Format categories.
3 CFPO: Content-Format Integrated Prompt Optimization
As we have established, the effectiveness of LLMs is profoundly influenced by both the content and format of prompts. Existing automated prompt optimization methods have largely overlooked the format dimension, which exhibits a strong model bias. To address this critical limitation, we introduce Content-Format Integrated Prompt Optimization (CFPO) framework that jointly optimizes both prompt content and format. This contrasts with prior approaches focusing solely on content optimization. Our goal is to identify an optimal prompt $p^{*}$, comprising both content ($c^{*}$) and format ($f^{*}$), that maximizes performance on an evaluation dataset $\mathcal{D}$, given by:
$p^{*}:(c^{*},f^{*})=\arg\max_{c\in\mathcal{L},f\in\mathcal{F}}m(c,f|\mathcal{D}),$ (1)
within the coherent natural language space $\mathcal{L}$ and the space of all possible formats $\mathcal{F}$, guided by a metric function $m(\cdot)$ that assesses the prompt’s quality.
To effectively search this complex space, CFPO employs a two-pronged iterative approach, detailed in Figure 2, that consists of two concurrently-run optimizers: a Component-wise Content Optimizer and a Format Optimizer. The content optimizer refines the textual content of a prompt, while the format optimizer explores the structural arrangement of its elements. Importantly, our framework acknowledges the inherent interdependence of content and format and thus iterates between optimizing them, to find their optimal combination. The following sections detail our structured prompt template (Section 3.1), our innovative format optimization approach (Section 3.2), and finally our integrated optimization process (Section 3.3).
3.1 Structured Prompt Template
To enable fine-grained and targeted optimization, our framework adopts a structured prompt template inspired by guidelines from  OpenAI (2024b) and Google (2024). This template decomposes prompts into distinct functional components, facilitating both analysis and selective mutations. Specifically, our template divides a prompt into content-based components and format-based components, as illustrated in Figure 3.
The Content-based Components are:
Task Instruction defines the primary goal, guiding the model’s overall behavior.
Task Detail offers supplementary task-specific information, including resolution steps.
Output Format specifies the desired output structure (e.g., JSON, bullet points, etc.).
Few-shot Examples provide contextual learning patterns, consisting of:
  • Examples: specific instances pertinent to the task, including inputs and expected outputs.
  • Example Hinter (optional): a brief hint indicating that examples segment will follow, e.g., ’Here are some examples:’.
  • CoT Hinter (optional): encourages a chain-of-thought reasoning process, e.g., ’Let’s think step by step’.
Query shows the question or request to be answered by the LLM.
The Format-based Components are:
Query Format: defines how to structure the rendering of examples and queries.
Prompt Renderer defines how to aggregate all components into a structured prompt.
The formulation of the structured prompt template is fundamental to our optimization approach. This design yields two key advantages: first, it facilitates a structured component functionality where each part serves a specific purpose, promoting a more organized prompting framework; second, it enables fine-grained optimization by decoupling format from content, thus allowing targeted and precise modifications of individual components.
3.2 Format Optimizer Design
The key aspect of our work is the format optimization methodology. To efficiently explore the extensive range of prompt formats, the CFPO format optimizer adopts an approach that utilizes a format pool with a scoring system and an LLM-assisted format generation module. It strategically explores, evaluates, and refines formatting choices, all while learning from previous iterations.
3.2.1 Format Pool with Scoring System
The format pool is designed to hold the format configurations we use to generate prompts. As shown in Figure 4, these configurations are separated into two dimensions: the Prompt Renderer, which dictates the overall structure of the prompt, and the Query Format, which governs the rendering of in-context examples and queries. This distinction allows us to explore both macro and micro-level formatting variations.
To dynamically evaluate the potential of each format, we developed a scoring system for assessing the performance of each format $f$, represented as $Q(f)$. This system updates the performance score of $f$ across various prompt contents using the formula $Q(f)\leftarrow Q(f)+\sum_{c}m(c,f)$, where $c$ represents each content instance in current round. Additionally, we maintain $N(f)$ to count the number of times a format has been visited, which facilitates score normalization.
To initialize the exploration, we constructed a initial format search space $\mathcal{F}$, comprising a set of predefined commonly used formats, as illustrated in Figure 4. We also incorporate diverse variations of these predefined formats into the initial search space, such as adjustments to spacing, punctuation, and the use of special symbols. This establishes a starting point for our optimization.
3.2.2 LLM-assisted Format Generation
The variability of format space requires an automated process for effective expansion and exploration. To that end, we introduce an LLM-based format generator, $LLM_{f\_gen}$, which autonomously generates new formats based on information in the existing format pool.
This evolutionary approach integrates the format generation into each optimization round, allowing for the creation of new and potentially beneficial formats. To enhance the efficiency of this process, we guide the LLM towards more promising areas by informing it of the performance function, $\frac{Q(f)}{N(f)}$. This iterative process not only diversifies the format pool but also ensures that our system can adapt to and incorporate a wide range of formats, thereby enhancing its utility and effectiveness. More detailed information of our format generation process is provided in the Appendix A.2.
3.2.3 Search Format via Format Optimizer
For each content candidate generated by the content optimizer, the format optimizer aims to identify the most appropriate format from format pool. To navigate the balance between exploring new formats and exploiting known effective ones, we implemented the Upper Confidence Bounds applied to Trees (UCT) algorithm (Kocsis and Szepesvári, 2006). The UCT algorithm employs a selection criterion given by:
$UCT(f)=\frac{Q(f)}{N(f)}+\alpha\sqrt{\frac{\sum_{f}N(f)}{N(f)}}$ (2)
where $\alpha$ serves as a balancing hyper-parameter, adjusting the trade-off between exploration and exploitation.
The overall process, outlined in Algorithm 1, selects $2k$ formats for evaluation in each optimization round: $k$ promising formats from the pool (based on UCT score), and $k$ new formats generated by the $LLM_{f\_gen}$. The selected formats from both the existing pool ($\mathcal{F}_{select}$) and the new generated pool ($\mathcal{F}_{gen}$) are then evaluated using a predefined metric function $m(\cdot)$, and the best-performing format among the tested candidates will be identified. The result is then incorporated into the pool for future iterations.
Algorithm 1 Searching Optimal Format Given a Prompt Candidate
0:  $p_{0}=(c_{0},f_{0})$: initial prompt, $p=(c,\cdot)$: current prompt candidate(with content $c$), $\mathcal{F}$: dynamic format pool, $k$: number of formats, $m(\cdot)$: evaluation metric, $\mathcal{D}$: evaluation data.
1:  Initialize: $Q(f)\leftarrow m(c_{0},f)$, $N(f)\leftarrow 1$ for all $f\in\mathcal{F}$
2:  Format Selection: $\mathcal{F}_{select}\leftarrow\{f\in\mathcal{F}:f$ is in the top $k$ w.r.t. $UCT(f)\}$
3:  Format Generation:
4:  for each $i=0,1,...,k$ do
5:     Generate format: $f_{new}\leftarrow LLM_{f\_gen}(\mathcal{F})$
6:     Collect $f_{new}$ to $\mathcal{F}_{gen}$, and add $f_{new}$ to $\mathcal{F}$
7:  end for
8:  Format Evaluation:
9:  for each $f\in\mathcal{F}_{select}\cup\mathcal{F}_{gen}$ do
10:     Evaluate $m(c,f)$ with dataset $\mathcal{D}$
11:     $Q(f)\leftarrow Q(f)+m(c,f)$
12:     $N(f)\leftarrow N(f)+1$
13:     Update $UCT(f)$ by Eq. 2
14:  end for
15:  $\hat{f}\leftarrow\arg\max_{f\in\mathcal{F}_{select}\cup\mathcal{F}_{gen}}m(c,f)$
15:  The optimal format $\hat{f}$ for content $c$
By iteratively evaluating formats, the format optimizer ensures a balance between exploring new formats and refining current ones, converging to the best format configuration.
3.3 Integrated Optimizer Design
CFPO orchestrates the Component-wise Content Optimization and Format Optimization within an iterative framework to jointly optimize content and format. This iterative process (illustrated in Figure 2) is key to our methodology.
Component-wise Content Optimization: This stage employs two primary strategies for mutating the content of prompts. The first strategy is case-diagnosis and revision, leveraging test cases to assess the efficacy of the current prompt. The outcomes of these test cases, including both correct and incorrect samples, are analyzed by the LLM optimizer. This optimizer evaluates the performance and pinpoints specific components in need of optimization. Subsequently, targeted feedback is applied to these identified components for enhancement, resulting in improved prompts. For example, if the output is not in the specified format, the output format component will be altered. Additionally, a Mote-Carlo sampling strategy is employed to enhance the optimization robustness by generating synthetic content with same semantics for randomly selected components. After this step, we select top-performing content candidates based on an evaluation dataset for the next stage.
Format Optimization: As discussed in Section 3.2, this stage identifies the most effective format for each candidate prompt content from the previous content-optimization step. The format optimizer applies our dynamic format exploration and evaluation process, tracking performance, and updating the format pool’s scoring system. The Format Optimizer meticulously tracks the performance of all evaluated formats, providing valuable insights to guide the selection of formats in subsequent iterations. Simultaneously, it retains only the most effective format for each prompt, ensuring the diversity of prompt content candidates during beam search.
In summary, the two optimizers work in tandem, leveraging the strengths of the LLM to facilitate swift adaptation and customization. Importantly, this iterative process allows for the optimization of format and content, thereby significantly enhancing the quality of the generated prompts.
Method Mistral-7B-v0.1 LLaMA-3.1-8B LLaMA-3-8B-Instruct Phi-3-Mini-Instruct
GSM8K
Baseline (1-shot cot) 36.85 50.03 74.00 83.45
Baseline (8-shot cot) 38.21 51.02 73.46 85.75
GRIPS 39.04 50.27 74.53 83.47
APE 40.33 52.39 75.13 83.85
ProTeGi 45.72 54.74 75.36 84.84
SAMMO 43.82 54.74 75.89 84.76
CFPO (Ours) 53.22 63.38 80.74 89.16
MATH-500
Baseline (1-shot cot) 4.60 10.58 12.20 12.60
Baseline (4-shot cot) 10.20 23.40 14.00 40.40
GRIPS 13.40 15.80 23.60 10.80
APE 11.60 12.80 22.80 30.60
ProTeGi 10.80 17.00 18.40 28.80
SAMMO 12.20 15.40 25.80 42.40
CFPO (Ours) 14.80 26.99 33.33 44.20
ARC-Challenge
Baseline 67.15 73.81 75.94 84.39
GRIPS 77.05 77.90 79.61 87.46
APE 75.85 77.05 78.67 87.63
ProTeGi 76.54 77.22 79.86 87.54
SAMMO 77.22 77.13 79.86 87.03
CFPO (Ours) 79.35 78.50 80.63 88.23
Big-Bench Classification
Baseline 56.00 64.00 70.00 54.00
GRIPS 86.00 67.00 84.00 69.00
APE 73.00 65.00 60.00 63.00
ProTeGi 83.00 81.00 82.00 76.00
SAMMO 86.00 80.00 86.00 78.00
CFPO (Ours) 94.00 90.00 91.00 87.00
Table 1: Main results on math reasoning tasks and commonsense reasoning tasks.
4 Experiments
4.1 Experimental Setups
Dataset and Models. To rigorously evaluate CFPO, we selected a diverse set of tasks and models. Our benchmark tasks span various domains and complexities, including:
  • Reasoning: GSM8K (Cobbe et al., 2021) and MATH500 (Hendrycks et al., 2021; Lightman et al., 2023) which require complex mathematical reasoning abilities.
  • Multiple-choice: ARC-Challenge (Clark et al., 2018), demanding understanding and selection among alternatives.
  • Classification: The Implicatures task from the Big-Bench benchmark (bench authors, 2023) to evaluate classification proficiency.
Our model selection includes a mix of foundational and instruction-tuned models to understand the generalizability of our approach:
  • Foundational Models: Mistral-7B-v0.1 (Jiang et al., 2023) and LLaMA-3.1-8B (Meta, 2024b) represent pre-trained models.
  • Instruction-Tuned Models: LLaMA-3-8B-Instruct (Meta, 2024a) and Phi-3-Mini-Instruct (Microsoft, 2024) represent models specifically fine-tuned for instruction following.
Furthermore, we use GPT-4 (2024-05-01-preview) as the LLM optimizer for content mutation and format generation (OpenAI, 2024a).
Implementation Details. The training process involved $20$ iterative rounds, each consisting of content and format optimization. During content optimization, case-diagnosis and Monte Carlo sampling each generate 4 prompts per round. A set of 40 test cases is used, with 5 correct and incorrect cases leveraged for case-diagnosis. The number of prompt-structured components decreases progressively from 4 to 1, narrowing the search space over time to enhance efficiency. For format optimization, 4 UCT-selected formats and 4 newly generated formats are used to generate new prompts. The coefficient in the UCT selection process $\alpha$ is set to $31e-$ 3. Beam search, with a budget of 8, is employed during mutations to ensure effective exploration. Eval data sizes are configured as 50, 300, 500, and 500 for BigBench-Classification, MATH500, GSM8K, and ARC-Challenge, respectively. The best performing prompt on the evaluation set for each method was selected and reported on the test set.
Baselines. To evaluate the effectiveness of CFPO, we compared againt several commonly used and popular baselines. GrIPS (Prasad et al., 2023) performs syntactic phrase-level edits in instruction, representing a non-LLM-based optimization approach. APE (Zhou et al., 2023) and ProTeGi (Pryzant et al., 2023) both employ LLM to optimize prompt content, but differ in mutation strategy. APE adopts an instruction induction approach, while ProTeGi leverages test cases feedback with LLM to guide the mutation process. SAMMO (Schnabel and Neville, 2024) introduces a structured framework that incorporates a preliminary format mutation strategy, which relies on random selection from a predefined format pool. This choice of baselines enables a comprehensive assessment of CFPO’s capabilities against various types of optimization approaches. All methods were evaluated using consistent experimental configurations to ensure a fair comparison.
Initial Prompts. To establish a reasonable starting point, we employed a single in-context example without any further instruction as the initial prompt for each model and task, except for GrIPS which requires an initial instruction. Chain-of-Thought examples were employed for the reasoning tasks. We also report common baseline prompts, including 8-shot for GSM8K and 4-shot for MATH500. A comprehensive list of our initial prompts is in Appendix C.
4.2 Main Results
Table 1 summarizes the performance of CFPO in comparison with several state-of-the-art methods across four datasets. The results highlight the superior performance of CFPO, significantly outperforming the baseline prompt as well as competing methods. We observed that pre-trained models exhibit greater sensitivity to prompt formatting, leading to substantial improvements when optimized by CFPO. Notably, optimized prompts for pre-trained models tend to be longer and incorporate more in-context examples, suggesting that these characteristics better align with the optimization needs of pre-trained models (see Appendix D.1). In contrast, instruction-tuned models display relatively more robust results and smaller gains, likely due to their inherent adaptability and generalization.
For the reasoning tasks, GSM8K and MATH, prompt optimization is especially impactful due to the sensitivity of these tasks to prompt structure. CFPO, which integrates unified content and format optimization, delivers significant performance gains. Specifically, the improvement for GSM8K is more evident compared to the more challenging MATH task, where the inherent complexity limits the magnitude of improvement. Moreover, feedback-based methods like ProTeGi, SAMMO, and CFPO, consistently outperform the other baselines because they leverage iterative feedback for prompt refinement. In contrast, GRIPS, which is limited to phrase-level mutations, exhibits marginal improvements. These results underline the effectiveness of the integrated optimization strategy adopted by CFPO. The selected optimal prompts discovered by our approach can be found in Appendix D.
4.3 Ablation Study
Impact of the Format Optimizer. CFPO incorporates a unique format optimization process, leveraging LLM for format generation and a UCT-based strategy for format selection. To evaluate its effectiveness, we evaluated two variations of our method: (1) CFPOc, which optimizes content while keeping format fixed, and (2) CFPOc+Format, which first optimizes content, then performs a separate format optimization step. Table 2 shows that both CFPOc and CFPOc+Format underperform compared to the full CFPO approach, highlighting the importance of the integrated content and format optimization approach. The need for a joint optimization process which addresses the interdependence of content and format is essential for prompt optimization.
Task Method LLaMA-3.1-8B LLaMA-3-8B-Instruct
GSM8K ProTeGi 54.74 75.36
CFPOc 58.07 77.71
CFPOc+Format 61.94 79.30
CFPO 63.38 80.74
BBC ProTeGi 81.00 82.00
CFPOc 85.00 85.00
CFPOc+Format 88.00 89.00
CFPO 90.00 91.00
[…]
Effectiveness of Format Generation. We compared the full CFPO approach against a variant that uses format from initial format pool without using LLM for generation. As presented in Table 4, CFPO with format generation consistently outperforms the baseline relying solely on the initial pool. These results demonstrate the effectiveness of the proposed format exploration mechanism in enhancing both the quality and diversity of prompts.
Task Method LLaMA-3.1-8B LLaMA-3-8B-Instruct
GSM8K w/o Format Gen 62.70 78.85
with Format Gen 63.38 80.74
BBC w/o Format Gen 88.00 87.00
with Format Gen 90.00 91.00
Table 3: Impact of format generation during prompt optimization.
Effectiveness of Format Selection. We further evaluated our UCT-based format selection process, compared it to a random selection from the format pool and a greedy selection without exploration (using $\alpha=0$ in Eq. (2)). As presented in Table3, CFPO consistently achieves the best performance across all experimental settings, demonstrating the efficacy of the UCT-based selection strategy.
Task Method LLaMA-3.1-8B LLaMA-3-8B-Instruct
Random 62.40 78.82
GSM8K UCT($\alpha=0$) 63.23 79.08
UCT(ours) 63.38 80.74
Random 85.00 87.00
BBH UCT($\alpha=0$) 86.00 88.00
UCT(ours) 90.00 91.00
Table 4: Impact of different format selection strategies during optimization.
Effectiveness of the Content Optimizer. As presented in Table 2, we include ProTeGi (Pryzant et al., 2023), a baseline that optimizes only the content. In contrast, our CFPOc, which incorporates structured prompting and integrates correct cases for diagnosis, achieves significant performance improvements, which highlights the effectiveness of our content optimization strategy.
5 Conclusion
This paper introduces Content-Format Integrated Prompt Optimization (CFPO), a innovative methodology that concurrently optimizes both prompt content and format. CFPO incorporates the Prompt Renderer and the Query Format within a structured prompt template. By leveraging distinct optimization strategies, CFPO discovers high-performing prompts that outperform content-only methods, addressing a critical gap in existing research. Our results demonstrate the substantial significant influence of format on LLM performance, underscoring the necessity of a joint optimization approach. These findings emphasize the importance of integrating content and format considerations in prompt engineering. CFPO represents a significant advancement, empowering developers to design effective and robust prompts and unlocking the full potential of LLMs across diverse applications.
Appendix A Appendix: Detailed Optimization Process and Meta-Prompts
A.1 Meta-Prompt Header Setup
At the beginning of the prompt, we introduce the task and provide a detailed explanation of the prompt’s components, followed by the current version of the prompt. Below is the structure of the meta-prompt header, where placeholders are denoted in [ALL CAPS]:
Im trying to write a prompt to [TASK INTENTION].
The current prompt consists of several key components, including:
[DESCRIPTION OF COMPONENTS]
The complete prompt is as follows:
"""[CURRENT PROMPT]"""
A.2 Format Generation
Our format generation process is a two-step procedure designed to create diverse and effective prompt formats. We focus on generating two key components of a prompt’s format: the Prompt Renderer and the Query Format. The appendix presents examples of the format generated using this pipeline.
Step 1: Format Description Generation. For each component (i.e., Prompt Renderer and the Query Format), we first generate a natural language description of the format, alongside an example of how this format would render a sample input. This description acts as a blueprint, guiding the subsequent code generation. We utilize a meta-prompt to instruct an LLM to perform this task. The meta-prompt takes existing format examples as context and generates new format descriptions along with rendered results. As an illustrative example, here is a conceptual outline of the meta-prompt employed for generating new Query Format descriptions:
[META PROMPT HEADER]
We have some preset QUERY_FORMAT candidates, here are our whole search pool:
[ALL EXISTING QUERY FORMATS DESCRIPTION]
Here are two examples from our QUERY_FORMAT candidates as for your reference:
<Format name: Question-Answer>
[RENDERED EXAMPLE 1]
<Format name: Instruction-Response>
[RENDERED EXAMPLE 2]
Please generate ONE new format for the QUERY_FORMAT segment, its description and render the provided example using this new format. The new format could either be a completely new format or a variation of an existing format.
If you choose to generate a completely new format, please ensure that the new format is conventional, structured, and aligned with commonly used query formats. Avoid overly creative or unconventional formats that deviate significantly from standard practices. The new format should be distinct from the existing formats.
The variation can focus on two parts, CASING and SEPARATOR:
CASING refers to both the capitalization of the text (e.g., f(x) = x.title(), f(x) = x.upper(), f(x) = x.lower()) and the specific wording or phrasing used (e.g., changing "question" to "instruction" or "input").
[…]
Note that focus solely on the format itself without altering the content of the question and answer. The format should remain focused on the existing structure (e.g., Question/Answer or Instruction/Response) without modifying the content or introducing any new sections. Avoid the use of underlines or any unconventional formatting styles among words. The format name should only include alphanumeric characters and underscores. Special characters such as ‘|‘, ‘!‘, ‘#‘, @‘, and spaces should be avoided.
Please encapsulate the new query format using the following format:
<START>
<Format name: [format name]>
<Description: [format description]>
[The example rendered by the newly generated format]
<END>
Step 2: Format Code Generation. Based on the natural language description and rendered example produced in Step 1, we subsequently generate the corresponding code implementation of the new format. This code will be used by the system to render prompts according to the defined format. We again leverage a meta-prompt to instruct the LLM, this time to generate the executable code. As an illustrative example, here is a conceptual outline of the meta-prompt employed for generating the code representation of a new Query Format:
[META PROMPT HEADER]
We have some preset QUERY_FORMAT candidates, here are our whole search pool:
[ALL EXISTING QUERY FORMATS DESCRIPTION]
Here are two code implementations from our QUERY_FORMAT candidates as for your reference:
<Format name: Question-Answer>
<Renderer code>
[Question-Answer RENDERER CODE]
<Extractor code>
[Question-Answer EXTRACTOR CODE]
<Format name: Instruction-Response>
<Renderer code>
[Instruction-Response RENDERER CODE]
<Extractor code>
[Instruction-Response EXTRACTOR CODE]
Here is the example rendered by the new format:
[RENDERED RESULTS]
[…]
Please encapsulate the code using the following format:
<START>
<Format name: {format_name}>
<Description: {format_description}>
<Renderer code>
[Renderer code]
<Extractor code>
[Extractor code]
<END>
A.3 Content Optimization
A.3.1 Case-diagnosis and Revision
As described in Section 3.3, content optimization is achieved through an iterative process of case-diagnosis and feedback guided mutation. To facilitate this process, we utilize three distinct meta-prompts, each tailored to a specific task within content optimization.
Case Diagnosis Meta-Prompt. This meta-prompt analyzes the current prompt’s performance against a set of test cases. It identifies areas for improvement and suggests specific modifications for the next iteration.
[META PROMPT HEADER]
Upon evaluating the current prompt, this prompt gets the following examples wrong:
[INCORRECT CASES]
Meanwhile, this prompt gets the following examples correct:
[CORRECT CASES]
Please review the provided examples of correct and incorrect answers, and identify [NUM OF DIAGNOSED COMPONENTS] specific area for improvement in the prompts. Each suggestion should focus on A SPECIFIC segment of the prompt that needs optimization. For each suggestion, provide a comprehensive explanation that encapsulates all the evaluation results. If you believe the EXAMPLES segment needs improvement, you may suggest one example that can be added, removed, or altered to enhance the EXAMPLES segment based on the examples given. If you think there is no need for improvement, do not return any prompt segment.
Please encapsulate each suggestion using the following format:
<START>
<Prompt segment: [Segment name]>
[Suggestion goes here]
<END>
Feedback Application Meta-Prompt. Based on the diagnosis, this meta-prompt generates targeted textual changes to enhance the prompt’s performance. It directly modifies the identified components of the prompt based on the feedback.
[META PROMPT HEADER]
The existing [COMPONENT NAME] segment contains:
[CURRENT CONTENT FOR THE COMPONENT]
Here are some suggestions for improving the [COMPONENT NAME] segments:
[GENERATED DIAGNOSES]
Based on the above information, I wrote [NUMBER OF GENERATED CONTENT] distinct and improved versions of the [COMPONENT NAME] segment within the prompt.
Each revised segment is encapsulated between <START> and <END>. In case this segment is an empty string, generate a suitable one referring to the suggestion.
The [NUMBER OF GENERATED CONTENT] revised [COMPONENT NAME] segments are:
Feedback Application Meta-Prompt (for Examples). This meta-prompt specifically handles the optimization of few-shot examples. It revises examples by adding, deleting, or modifying one single instances, ensuring that the in-context learning process is effective.
[META PROMPT HEADER]
The existing EXAMPLES segment contains:
[CURRENT IN-CONTEXT EXAMPELS IN PROMPT]
Here are some suggestions for enhancing the EXAMPLES segment:
[GENERATED DIAGNOSES]
Based on the above information, I have crafted [NUMBER OF GENERATED EXAMPLES] improved version of the EXAMPLES segment within the prompt. Each revision represents ONLY ONE of the following specific actions:
1. Addition: Incorporating one new example into the existing set.
2. Deletion: Eliminating one single example from the current set.
3. Modification: Changing the content of an example while maintaining its contextual relevance.
Please present the results without indicating which action was taken. Each refined EXAMPLES segment is marked by <START> and <END>.
The [NUMBER OF GENERATED EXAMPLES] revised EXAMPLES are:
A.3.2 Monte-Carlo Sampling
Monte-Carlo Sampling Meta-Prompt explores a wider range of semantically equivalent yet syntactically varied instructions, enhancing the chances of discovering more effective prompts.
[META PROMPT HEADER]
Please create a different version of [COMPONENT NAME] segment without changing its semantic meaning. In case this segment is an empty string, generate a suitable one. The existing [COMPONENT NAME] segment contains:
[CURRENT CONTENT FOR THE COMPONENT]
The varied [COMPONENT NAME] segment is as follows:
Monte-Carlo Sampling Meta-Prompt (for Examples) refines few-shot examples by strategically adding, deleting, or modifying single instances to ensure their effectiveness.
[META PROMPT HEADER]
The existing EXAMPLE set contains:
[CURRENT IN-CONTEXT EXAMPELS IN PROMPT]
Please generate a variation of the EXAMPLES set within the prompt while keeping the semantic meaning. The revision shoud represent ONLY ONE of the following specific actions:
1. Addition: Incorporating one new example into the existing set.
2. Deletion: Eliminating one single example from the current set.
3. Modification: Changing the content of an example while maintaining its contextual relevance.
Please present the results without indicating which action was taken. The varied EXAMPLES segment is as follows:
Appendix B Appendix: Examples of Generated Format
Here we select several format generated by GPT4 in CFPO process.
B.1 Query Format
QA_Titlecase_Separator
Question || In 3 years, Jayden will be half of Ernestos age. If Ernesto is 11 years old, how many years old is Jayden now?
Answer || Lets think step by step. Ernesto = 11 + 3 = <<11+3=14>>14 Jayden = 14/2 = <<14/2=7>>7 in 3 years Now = 7 - 3 = <<7-3=4>>4 Jayden is 4 years old.
QA_Brackets_Colon_Newline
[Question]:
In 3 years, Jayden will be half of Ernestos age. If Ernesto is 11 years old, how many years old is Jayden now?
[Answer]:
Lets think step by step.
Ernesto = 11 + 3 = <<11+3=14>>14 Jayden = 14/2 = <<14/2=7>>7 in 3 years Now = 7 - 3 = <<7-3=4>>4 Jayden is 4 years old.
QA_CapsBold_ColonNewline
**QUESTION**:
In 3 years, Jayden will be half of Ernestos age. If Ernesto is 11 years old, how many years old is Jayden now?
**ANSWER**:
Lets think step by step.
Ernesto = 11 + 3 = <<11+3=14>>14 Jayden = 14/2 = <<14/2=7>>7 in 3 years Now = 7 - 3 = <<7-3=4>>4 Jayden is 4 years old.
Cascading_Statements
Question: Statement 1 | Every element of a group generates a cyclic subgroup of the group. Statement 2 | The symmetric group S_10 has 10 elements.
Options:
-A True, True
-B False, False
-C True, False
-D False, True
Answer: C
Highlight_Separator_Case
QUESTION > Statement 1 | Every element of a group generates a cyclic subgroup of the group. Statement 2 | The symmetric group S_10 has 10 elements.
OPTIONS > (A) True, True (B) False, False (C) True, False (D) False, True
ANSWER > C
B.2 Prompt Renderer
Concise_Bullet_Points_Renderer
- Task Instruction: Write a function that returns the sum of two numbers.
- Task Detail: The function should take two numbers as input and return their sum.
- Examples: Input: 1, 2
Output: 3
- Query: Input: 1, 2
Output:
Tabular_Sections_Renderer
| Task Instruction | Write a function that returns the sum of two numbers. |
| Task Detail | The function should take two numbers as input and return their sum. |
| Examples | Input: 1, 2
Output: 3 |
| Query | Input: 1, 2
Output: |
Checklist_Format_Renderer
- [ ] **Task Instruction**
Write a function that returns the sum of two numbers.
- [ ] **Task Detail**
The function should take two numbers as input and return their sum.
- [ ] **Examples**
Input: 1, 2
Output: 3
- [ ] **Query**
Input: 1, 2
Output:
Appendix C Appendix: Initial Prompt
C.1 GSM8K
Prompt Renderer: Directly Joint Query Format: QA
Q: There are 15 trees in the grove. Grove workers will plant trees in the grove today. After they are done, there will be 21 trees. How many trees did the grove workers plant today?
A: There are 15 trees originally. Then there were 21 trees after some more were planted. So there must have been 21 - 15 = 6. The answer is 6.
{{Query placeholder}}
C.2 MATH500
Prompt Renderer: Directly Joint
Query Format: Question-Answer
A chat between a curious user and an AI assistant. The assistant gives step-by-step solutions to the users questions. In the end of assistants response, a final answer is given in the format of "The answer is: <ANSWER>.".
Here are some examples:
[…]
{{Query placeholder}}
C.3 ARC-Challenge
Prompt Renderer: Directly Joint
Query Format: MultiChoice_QA
You are a commonsense helper. I will provide several examples and a presented question. Your goal is to pick the most reasonable answer among the given options for the current question. Please respond with the corresponding label (A/B/C/D) for the correct answer.
Here are some examples:
Question: Forests have been cut and burned so that the land can be used to raise crops. Which consequence does this activity have on the atmosphere of Earth?
Choices:
A: It reduces the amount of carbon dioxide production
B: It reduces the production of oxygen
C: It decreases the greenhouse effect
D: It decreases pollutants in the air
Answer: B
{{Query placeholder}}
C.4 Big-Bench Classification
Prompt Renderer: Directly Joint
Query Format: Input-Output
Examples:
Input: Speaker 1: You do this often?’ Speaker 2: Its my first time.’
Output: no
{{Query placeholder}}
Appendix D Appendix: CFPO Results Analysis
D.1 In-context Examples and Text Length
Figure 5: Overview of in-context examples and text lengths for various tasks and models.
Figure 5 presents an overview of the number of in-context examples and the text length of optimized prompts across various tasks and models. An interesting pattern emerges: pre-trained models consistently prefer prompts with longer text and more in-context examples compared to instruction-tuned models. This observation suggests that pre-trained models benefit more from explicit context and detailed reasoning steps, which align with their less task-specialized nature. In contrast, the relative insensitivity of instruction-tuned models to prompt length and in-context examples supports the notion that these models have already trained with task-specific knowledge during fine-tuning, reducing their dependence on highly detailed prompts.
D.2 Examples of Optimal Prompt
Here we selected several optimal prompts searched by CFPO.
LLaMA-3.1-8B on GSM8K
**Understanding the Task: A Foundation for Mathematical Problem-Solving**
Your task is to methodically analyze the information provided and logically deduce the correct answer to the mathematical problem. Delve into each relevant detail, ensuring no critical step or aspect is overlooked. Approach the solution with a detailed-oriented mindset, ensuring every part of the process is considered to arrive at an accurate conclusion. Reflect on all the elements that might influence your reasoning or calculation, striving for thoroughness in your analysis.
**Decoding Mathematical Language in Real-World Scenarios**
For the most effective problem-solving in mathematics, particularly when faced with intricate calculations over periods or under specific scenarios affecting results, an attentive and systematic method is key. Start by accurately determining the base numerical value. Then proceed by methodically listing every significant change whether it be increases, decreases, or modifications that impacts this base figure as the scenario unfolds, making sure to include each change in your overall computations. Its essential to focus on the concept of compounded operations, whether theyre applied annually, monthly, or daily, and to thoughtfully evaluate the consequences of extraordinary events or circumstances (like an unexpected inheritance, a yearly loss, or a singular occurrence with a major impact) that might significantly shift the end calculations. Sharpen your attention on the dynamics of numerical relationships, particularly in cases involving ratios, proportions, and the impact of percentage changes over durations, to avoid common mistakes. Misunderstandings or misapplications of these numerical relationships can frequently cause inaccuracies. Thus, it is critical to scrutinize these mathematical relationships, whether they are of direct or inverse proportions, as well as the aggregate effects of consecutive percentage changes, as outlined in the problem description. This intensified attention is pivotal for an accurate and detailed resolution of complex issues, marked by multiplicative elements and interconnected circumstances. Reflect deeply on the significance of every step in the calculation process, absorbing the nuances of these changes, to systematically arrive at the most precise solution.
**Ensuring Your Solution Fits the Scenario Perfectly**
In presenting your solution, ensure it comprises both a numerical answer and a meticulously detailed explanation of the process leading to it. Begin with outlining the initial conditions and sequentially narrate the calculations you make at each step, highlighting any compounded operations or adjustments made to account for unique scenarios or conditions. This progression should clearly show how each step contributes to arriving at the final answer. For instance, if the task involves calculating the total costs saved over time with additional periodic benefits, your response should methodically explain: "Starting with an initial savings of X, plus Y every Z period, and considering an additional benefit of A every B period, leads to a total of...". This comprehensive breakdown not only bolsters the understanding of the mathematical principles applied but also provides a robust framework for identifying and rectifying any potential inaccuracies throughout the problem-solving process.
**Examples to Illuminate the Path**
To better grasp the concepts, consider the following illustrative examples:
Question: There are 15 trees in the grove. Grove workers will plant trees in the grove today. After they are done, there will be 21 trees. How many trees did the grove workers plant today? / ANSWER: Think through the problem step by step, diving into each segment for a thorough exploration to piece together the final answer. There are 15 trees originally. Then there were 21 trees after some more were planted. So there must have been 21 - 15 = 6. The answer is 6.
Question: A book club starts with a membership of 120. If the club increases its membership by 10% in the first year and then loses 5% of its members in the second year, what is the total membership at the end of the second year? / ANSWER: Think through the problem step by step, diving into each segment for a thorough exploration to piece together the final answer. The club starts with 120 members. In the first year, it increases by 10%, which is 0.10 * 120 = 12, so there are 120 + 12 = 132 members after the first year. In the second year, the club loses 5% of its members, which is 0.05 * 132 = 6.6, but since the number of members must be an integer, we consider a loss of 7 members (assuming the figure is rounded up for practical reasons). Therefore, there are 132 - 7 = 125 members at the end of the second year.
Question: Martin saves \$10 every week. In addition, every third week, he earns an extra \$15 from helping his neighbor. How much has Martin saved after 9 weeks? / ANSWER: Think through the problem step by step, diving into each segment for a thorough exploration to piece together the final answer. Martin saves \$10 each week, so over 9 weeks, he saves 9 * \$10 = \$90. Additionally, every third week, he earns an extra \$15, which occurs three times within 9 weeks (in the 3rd, 6th, and 9th weeks). So, he earns an extra 3 * \$15 = \$45 from helping his neighbor. Therefore, the total amount Martin has saved after 9 weeks is \$90 + \$45 = \$135.
Question: A teacher divides a class into groups for a project. If the ratio of boys to girls in the class is 3 to 2, and there are 30 students in the class, how many boys are in the class? / ANSWER: Think through the problem step by step, diving into each segment for a thorough exploration to piece together the final answer. The total ratio units for boys to girls in the class is 3 + 2 = 5. With 30 students in the class, each ratio unit represents 30 / 5 = 6 students. Therefore, the number of boys, represented by 3 parts of the ratio, is 3 * 6 = 18. The answer is 18.
Question: Grandma wants to order 5 personalized backpacks for each of her grandchildrens first days of school. The backpacks are 20% off of \$20.00, and having their names monogrammed on the backpack will cost \$12.00 each. How much will the backpacks cost in total? / ANSWER: Think through the problem step by step, diving into each segment for a thorough exploration to piece together the final answer. The backpacks are 20% off of \$20.00, so the price after the discount is \$20.00 - (\$20.00 * 20%) = \$20.00 - \$4.00 = \$16.00 each. The monogramming costs an additional \$12.00 per backpack. Therefore, the total cost for each backpack is \$16.00 + \$12.00 = \$28.00. For 5 backpacks, the total cost will be 5 * \$28.00 = \$140.00. The correct answer is \$140.00.
**Query**
{{query}}
LLaMA-3-8B-Instruct on MATH-500
- Task Instruction: A chat between a curious user and an AI assistant focused on solving mathematical and reasoning tasks. The assistant is expected to deliver step-by-step solutions to the users questions, emphasizing mathematical accuracy and rigor throughout the process. It must ensure that each mathematical operation and logical deduction is carefully examined and validated to derive the correct solution. At the conclusion of the response, the final answer should be presented in the format of "The answer is: <ANSWER>.", thereby confirming the solutions validity and demonstrating a thorough understanding of the problem-solving approach.
- Task Detail: In addressing equation-based inquiries, precision in algebra, geometry, piecewise functions, complex numbers, and financial mathematics is paramount. This involves a detailed analysis of each equation, assessing every element and specific condition. For piecewise functions, its critical to ensure continuity by solving for variables that maintain consistency across sections. In geometry, integrating measurements such as angles, lengths, and areas is fundamental. Algebraic queries require a consideration of all potential solutions and constraints, ensuring a comprehensive resolution. The addition of complex numbers into this mix necessitates a thorough understanding of their properties and operations to accurately determine both real and imaginary solutions. Similarly, tackling financial mathematics problems demands a deep comprehension of concepts such as compound interest, present value, and future value to make precise financial forecasts and comparisons. This holistic approach confirms that all aspects of the problem are considered and that the solution accounts for every requirement, assuring mathematical integrity in the resolution process.
- Output Format: 1. Solutions that involve fractions, square roots, or crucial mathematical figures (e.g., pi) must be simplified to their most fundamental form. This includes reducing fractions to their lowest terms and expressing square roots in their least complex radical form.
2. Avoid the use of decimals unless the question explicitly requires it or they are necessary for conveying the most precise value possible.
3. Present solutions involving square roots in their reduced radical form, ensuring the simplification process enhances comprehension without diluting mathematical integrity.
4. In scenarios involving complex numbers, represent answers in their standard form (a + bi), ensuring both a and b are presented in their simplest, most refined state. This emphasizes the need for a clear, coherent representation of solutions encompassing complex numbers.
5. Conclude your explanation with the statement: "The answer is: \[<ANSWER>\].", reinforcing consistency and clarity across various mathematical challenges. This concluding statement should encapsulate the solution in its simplest and most direct form, reflecting a thorough simplification and rationalization process.
Your explanation must delineate a detailed, step-by-step progression leading to the final solution. This approach is not merely about arriving at the correct answer but about illuminating the path taken to get there, ensuring a deep understanding and clear demonstration of the reasoning behind each step.
- Examples: Here are some examples:
### Instruction:
A rectangle ABCD has sides AB = 8 units and BC = 6 units. A circle with a radius r units is inscribed within this rectangle. Calculate the radius r of the inscribed circle, ensuring the answer is in its simplest form.
### Response:
Well approach this problem by breaking it down into manageable steps. We start by understanding that the radius of the inscribed circle is equal to the distance from the center of the rectangle to any of its sides because the circle is perfectly inscribed. In a rectangle, this distance is half the length of the rectangles shorter side. Therefore, the radius r of the inscribed circle is half the length of BC, which is \$6 $\div$ 2 = 3\$ units. The answer is: \$r=3\$.
### Instruction:
Given a triangle where two sides are represented by complex numbers (3 + 4i) units and (1 - 2i) units, and the angle between them is 90 degrees, calculate the length of the hypotenuse. Ensure your answer includes a comprehensive breakdown of complex number operations and geometric principles applied.
### Response:
[…]
- Query:
{{query}}
LLaMA-3.1-8B on ARC-C
<div class=’TaskInstruction’>
<h2>TaskInstruction</h2>
<p>Your mission is to meticulously assess each situation presented alongside a specific question, employing your critical thinking and analytical skills. Your task comprises not only identifying the most logical and coherent choice (A/B/C/D) but also thoroughly evaluating how each option connects or diverges from the questions essence. This requires a deep engagement with both the query and the choices, ensuring your reasoning is firmly anchored in the specifics of the options provided. It is essential to weave direct elements from the choices into your analysis, demonstrating a detailed understanding of how each option relates to the core question, and articulating why alternatives may be less fitting given the scenario. This approach ensures a nuanced and well-justified selection process, grounded in the interplay between the question context and the specific details of the available choices.</p>
</div>
<div class=’TaskDetail’>
<h2>TaskDetail</h2>
<p>In addressing the questions set before you, it is imperative to delve deeper than mere superficial observations or initial judgments. Each scenario or question must be examined not just in its immediate context but within a broader spectrum, looking into the underpinning mechanisms or far-reaching effects of each option presented. This necessitates a thorough exploration of the larger implications and the scientific or logical foundations that dictate the outcomes. For instance, in environmental matters, it is vital to assess not just the immediate effects but the sustained impact on the ecosystem. In the realm of science, such as when discerning chemical processes, it is crucial to understand the molecular or atomic level changes that classify a reaction as a chemical change. This enhanced level of scrutiny and deeper analysis will lead to more accurate and well-founded choices, ensuring your responses are not just correct, but are also backed by a solid understanding of the underlying principles or long-term consequences.</p>
</div>
<div class=’OutputFormat’>
<h2>OutputFormat</h2>
<p>For every query presented, your task is to identify the right choice from the options (A/B/C/D) accompanied by a concise rationale for your selection. This format is vital as it showcases the thought process leading to your decision, facilitating a comprehensive grasp and interaction with the task.</p>
</div>
<div class=’Examples’>
<h2>Examples</h2>
<p>Here are some examples:
Question: Forests have been cut and burned so that the land can be used to raise crops. Which consequence does this activity have on the atmosphere of Earth?
A: It reduces the amount of carbon dioxide in the atmosphere
B: It reduces the availability of oxygen
C: It lessens the greenhouse effect
D: It lowers the levels of pollutants in the air
Answer: B
Question: What is the most critical practice to ensure electrical safety while operating devices?
A: Ensure the device does not come into contact with water.
B: Use the device with hands covered in oil.
C: Operate the device with wet hands.
D: Leave the device plugged in when not in use.
Answer: A
Question: Placing a plant cell in a hypertonic solution typically results in which of the following?
A: The cell expanding as it absorbs water.
B: No significant change due to the rigid cell wall.
C: The cell shrinking as water exits the cell.
D: Rapid division of the cell.
Answer: C
Question: What is the primary effect of using fossil fuels on global climate change?
A: It leads to a significant reduction in greenhouse gases.
B: It decreases the Earths surface temperature.
C: It increases the amount of greenhouse gases in the atmosphere.
D: It contributes to a decrease in carbon dioxide levels.
Answer: C
Question: The process of photosynthesis in plants primarily involves which of the following transformations?
A: Converting oxygen and glucose into carbon dioxide and water
B: Transforming water and carbon dioxide into oxygen and glucose
C: Changing sunlight into chemical energy without producing oxygen
D: Producing carbon dioxide and glucose from oxygen and water
Answer: B
{{ query }}